In a catastrophic emergency on the space station,
Answer #1

speed of debris = 1.8 km/s

safe distance = 10 km

time taken to reach safe distance = 10 / 1.8 = 5.6 sec

consider the motion of pod :

launch speed , Vo = 50 m/s

distance of thrust firing point = 0.5 km

distance to be travelled to reach the safe point, d = 10 - 0.5 = 9.5 km = 9500 m

a = acceleration = 4 x 9.8 = 39.2 m/s2

using the equation :

d = Vo t + (0.5) a t2

9500 = 50 t + (0.5) (39.2) t2

t = 20.8 sec

so debris will hit the pods